OpenStudy (anonymous):

Please help for this elasticity problem! Demand Curve is given by q = (470 - p)^2/100 (0 <= p <= 470) where q is the number of copies the publisher can sell per week if it sets the price at $p. (a) Find the price elasticity of demand when the price is set at $31 per copy (b) Find the price at which the publisher should sell the books in order to maximize weekly revenue (c) What, to the nearest $1, is the maximum weekly revenue the publisher can realize from sales

OpenStudy (anonymous):

ahh business D:

OpenStudy (anonymous):

Yes, unfortunately! I really need help with this particular problem.

OpenStudy (dan815):

so set the p=31, and what do u get for q?

OpenStudy (anonymous):

1927.21?

OpenStudy (dan815):

kay so thats q, and q is how many copies he sells per week

OpenStudy (dan815):

so he gets to sell 1927.21 copies per week at $31 dollars a copy, how much money is that a week

OpenStudy (anonymous):

59743.51?

OpenStudy (dan815):

yep sounds about right

OpenStudy (dan815):

okay now for part b)

OpenStudy (dan815):

notice its not the q u want to maximise but, q*p

OpenStudy (dan815):

q is just the number you sell that week q*p is the revenue you get that week

OpenStudy (dan815):

q*p = (470 - p)^2/100 *p so maximize this function

OpenStudy (dan815):

Revenue = q*p = (470 - p)^2/100 *p Maximizing Revenue d/dp(Revenue)=d/dp( (470 - p)^2/100 *p)

OpenStudy (dan815):

hello?

OpenStudy (anonymous):

Sorry, for part a) the answer I put was not correct. :/

OpenStudy (dan815):

oh okay what does price elasticity mean

OpenStudy (dan815):

by the word itself im gonna assume how the price is changing with respect to demand

OpenStudy (dan815):

or maybe how the revenue is fluctuating around 31 dollars

OpenStudy (dan815):

dQ/dp @ p=31 =?

OpenStudy (dan815):

q = (470 - p)^2/100 dq/dp=-2(470-p)/100 dq/dt @ p=31 : -2(470-31)/100=-2*4.39=-8.78